OpenStudy (anonymous):

State how many imaginary and real zeros the function has. f(x) = x3 + 5x2 - 28x - 32

OpenStudy (callisto):

Use factor theorem to find the factor of f(x). By trial and error, we got f(-1) = 0 So, x+1 is one of the factors. Find the other factor(s) by division. Then, set f(x) = 0 x3 + 5x2 - 28x - 32 = 0 (x+1)(other factor(s)) = 0 x+1=0 or other factor(s) = 0. It shouldn't be difficult.